# Scaling Frustration Index and Corresponding Balanced State Discovery for Real Signed Graphs

### v2

@@ C.2 Implementation
The baseline implementation relies on the published Binary Linear Programming (BLP) code [Aref, 2021]. The binary linear model runs on a Jupyter Notebook in Python [Aref, 2021] and is based on a Gurobi mathematical solver and has several parameters like the termination parameter where we can set a time limit on how long the optimization process should last in Gurobi [Gurobi, 2008]. The binary terms in the objective function depend on the single AND constraints and two standard XOR constraints per edge, respectively. The graphBpp implementation extends the open-source implementation [Alabandi et al., 2021] to include and test proposed tree sampling strategies while keeping the original speedup optimization for finding fundamental cycles intact. The graphBpp algorithm works with different tree-sampling strategies. It is achieved by using C++ and involves minimizing the number of loops used, incorporating OpenMP directives for parallel processing, and promptly freeing up memory resources when they are no longer needed. In the implementation of Algorithm 5, the code checks the total RAM size of the Linux system during runtime and the amount of memory currently being used by the frustration cloud. These two values are compared in each iteration to decide how many balanced states can be stored. Two replacements in the ABS model’s objective function linearized two absolute value terms [Aref and Neal, 2021]. The code [Aref, 2021] was run with the following modifications: (1) -1 for the method parameter that indicates that the method for optimization is automatic, and the setting will typically choose the non-deterministic concurrent method in the Gurobi’s documentation [Gurobi, 2008] for this linear programming problem; (2) the lazy parameter is set to 1 with enabled speedup; (3) thread parameter is set to multiprocessing.cpu_count(, and (4) the time limit for the model run is set to up to 30 hours. Note that the value of the lazy attribute influences how aggressively the model is constrained. A value of 1 allows the constraint to cut off a feasible solution. The code provided [Aref, 2021] generates random graphs based on the specified number of nodes, edges, and probability of negative edges. Our improvements to the code allow for the code to (1) accept the same input format as graphBpp and to (2) detect and eliminate duplicates, inconsistencies, self-loops, and invalid signs in the input graph. The code for graphBpp is available on GitHub, and the data it uses is publicly accessible. The references for these resources are [Anonimoys, 2023] for the code and [Leskovec and Krevl, 2014, Kunegis, 2013, He and McAuley, 2016] for the data. On the other hand, for the gradient-based heuristic, we set $\lambda$ to 1000 and the learning rate $\alpha$ to 0.001.
@@ Appendix D Tree-Sampling techniques for graphBpp
The Depth-First Search (DFS) algorithm [Cormen, 2009] is a graph traversal method characterized by its time complexity of O($|V|+|E|$). The traversal commences at a root vertex and continues to explore as deeply as possible along each pathway (branch) before backtracking. This process repeats until it reaches the vertex where all adjacent vertices have already been visited. The Breadth first search (BFS) algorithm [Cormen, 2009] with time complexity O($|V|+|E|$) is a graph traversal approach in which the algorithm first passes through all vertices on the same level before moving on to the next level. A graph traversal technique is a strategy employed to visit all the vertices of a graph. A level is a group of vertices that are equidistant from the root vertex. We propose to use the randomized algorithms as follows: in each iteration, we shuffle and randomize a node’s neighborhood using a uniformly distributed random seed number before applying a static algorithm. The idea is that a vertex establishes a link to the first unvisited vertex based on the network’s randomized order of the adjacency list. Randomized Depth First Search (RDFS) algorithm transforms DFS into a non-deterministic algorithm by eliminating the static ordering of the adjacency lists. The time complexity of the DFS is known to be O($|V|+|E|$), where $|V|$ is the number of vertices and $|E|$ is the number of edges in the signed network. The algorithm also runs in linear time O($n$), where $n$ is the number of vertices adjacent to a specific vertex in the network, so the total time complexity is O($|V|+|E|$). Aldous-Broder algorithm with complexity O($|V|$) produces a random uniform spanning tree by performing a random walk on a finite graph with any initial vertex and stops after all vertices have been visited [Hu et al., 2021]. For the popular Kruskal’s algorithm [Wu et al., 2023] that has a time complexity O($|E|log|V|$) or O($|E|log|E|$), we intend to generate random spanning trees by assigning random weights to every edge in each iteration before running the algorithm. The method finds the minimum spanning tree of a connected and weighted graph. Randomizing the weights of Prim’s algorithm [Gass and Fu, 2013] [Wu et al., 2023] with complexity O($|V|^{2}$) can also generate random spanning trees.
@@ Appendix E Scaling graphBpp for Large Signed Graphs
The graphB+ algorithm [Alabandi et al., 2021] efficiently discovers fundamental cycles for the spanning tree $T$ and computes vertex and edge labels with linear time complexity. The algorithm requires a linear amount of memory, and the running time for balancing a cycle is linear in the length of the cycle times the vertex degrees but independent of the size of the graph. The labels here are specific values assigned to the vertices and edges after sampling a spanning tree as a preliminary step prior to finding and traversing the fundamental cycles and balancing them. These labels aid in finding these cycles and ensure efficient traversal. The algorithm speeds up the balancing to more than 14 million identified, traversed, and balanced fundamental cycles. Next, the algorithm traverses or visits each cycle in a specific order and takes note of which edge signs switched within each cycle to obtain an even number of negative edges along that cycle. The graphBpp, which builds on graphB+, scales the computation for large graphs as in Algorithm 5. First, we compute the number of balanced states that we can keep in the memory as $\mathcal{F}_{max}$ (Algorithm 5, line 3). Next, we keep only the $\mathcal{F}_{max}$ best-balanced states in $\mathcal{B}$, their count through all $k$ iterations in $\mathcal{C}$, and the number of switched edges for each of them in $\mathcal{S}(i)$. Note that there can be different balanced states of $\Sigma$ with the same number of switched edges. Finally, we compute the frustration index as a minimum of $\mathcal{S})$. The proposed approach scales well with the size of the signed graph, as we limit the number of the nearest balanced states that we keep in the memory based on their frustration index and capacity of the frustration cloud map in-memory storage, as illustrated in Algorithm 5. The comparison of balanced states now has up to $k$ iterations times $\mathcal{F}_{max}$ closest balanced states. We determine $\mathcal{F}_{max}$ so that the size of the frustration cloud in memory is smaller than $CAP$. In experiments, we define $CAP$ as 75% of the total RAM size, assuming that some vital system processes are running in the background. Note that $\mathcal{F}_{max}$ can be determined in the implementation by checking the current memory usage after every balanced state insertion and verifying that this insertion did not reach the limit. The benefit of saving balanced states in memory is that it allows us to record and study the frequency of each state and use them for other tasks, such as community detection and finding the largest balanced sub-graph.
@@
The size of the frustration cloud grows with the number of iterations as the probability of the previously unseen nearest balanced state grows. The size of the frustration cloud can become an issue for graphs with millions of vertices and vertices as the frustration cloud is too big for the main memory. The underlying data structure for the implementation in C++ is $std::map<std::string,std::pair<int,int>>$. The key, as discussed above, is of type string that represents a balanced state, and the pair stores two integers, one for the number of switches and the other for the frequency of the corresponding stable state. The string keys are indeed slower than using integer key types in the map data structure in regards to comparing keys. However, in our case, storing the balanced states using strings or integers should have a similar comparison performance, and the integer key approach would be much more complicated. First, if we use $std::map<int,std::pair<int,int>>$, then it would be difficult and more complex to represent the signed network using one integer key, whereas the string can intuitively concatenate all the edges with their signs to represent that graph. There is no direct way of representing this collection of edges forming the signed graphs using solely the integer key. If we were to use an integer as a key for the map, we would still have to loop through the edges in that key for the comparisons, and the cost would be $O(|E|)$ equivalent to using the string key.
@@
Our approach is different than the work by Tang et al. [Tang and Zhu, 2023] in several ways. First, we estimate the different types of latent variables, and we only use a random float vector $\Gamma$ of size equal to the number of vertices to determine the optimal membership of each vertex in the Harary subsets. Second, we are not modeling any signed networks. Our focus is to flip the sign of the edges after estimating the latent variables of each vertex and approximate the frustration index. Tang et al. did not explicitly and directly modify the edge signs of the graph. Third, we propose the usage of a loss function for computing the frustration index directly and for computing the gradients. Fourth, we use the vanilla gradient descent approach instead of the projected gradient descent to minimize our loss function. We simply threshold the latent variables after optimization (ex, if an element in the latent vector is 25 after the gradient-descent step, which is above 0, we assign that element to be 1). Fifth, our loss function is adopted from Du et al. [Du et al., 2016] in which they propose to measure imbalance for fully signed networks. The loss function used by Tang et al. is $L=\sum_{i<j}^{n}{|A_{ij}|}\frac{1+A_{ij}}{2}\eta_{ij}+|A_{ij}|\log(1-\Sigma(\eta_{ij})))$ where $\eta_{ij}=v_{i}v_{j}$, $A$ is the adjacency matrix, $\sigma$ is the sigmoid function, and $n$ is the number of nodes.
@@ G.2 The graphL Complexity Analysis
For every gradient update and computation, it is sufficient to loop over every edge in the signed graph in order to update the elements in the $\Gamma$ vector. The $\lambda$ is the number of gradient updates, and the total time complexity becomes O($\lambda*|E|$). We utilize the compressed sparse row (CSR) format to model the signed graph instead of the adjacency matrix because CSR scales better memory-wise. The construction of the adjacency matrix has a space complexity of O($V^{2}$), which isn’t computationally feasible when dealing with large signed graphs. Computing the amount of imbalance also occurs in constant time in each gradient update iteration, and it does not affect the total time complexity.
@@ Appendix H More Results on the Comparison between BLP and graphBpp
The execution time of BLP and graphBpp with all seven tree-sampling techniques is reported on the log 10 scale in seconds in Figure 5 (bottom) as BLP takes 30 hours for larger datasets (far right navy bar in Fig. 5 (bottom)). RDFS-BFS is competitive with BFS in terms of frustration index as a percentage of the total number of edges in the graph (green and blue bars in Figure 5 (top)) with the small timing overhead for large graphs (Fig. 5 (bottom)): BFS produces 117,587 frustrations while BFS-RDFS produces 115,932 frustrations in 1000 iterations for the slashdot dataset. The Prim approach is too slow for large datasets, and the baseline BLP takes too long or it does not complete. We also tabulated the timing per iteration for each tree-sampling technique in Table 8. Since the time complexity of Prim is O($V^{2}$), the number of iterations is set to 1, and it was very inefficient and slow for large graphs such as WikiConflict. Moreover, Aldous-Broder (AB) did not terminate for DutchCollege because, in uncommon scenarios. The AB approach can end up endlessly looping during random walks after the algorithm has visited all the current vertex’s neighbors.
@@ Appendix I Concluding Remarks
There is more than one way to achieve balance in the network. The frustration index characterizes the optimal nearest balanced state where the minimum edge switches are required to achieve balance in the network. The tree-spanning approach to graph balancing produces the nearest balanced states, e.g., there can be no other balanced state nearest balanced state derived from [Rusnak and Tešić, 2021]. In this paper, we extend our findings and propose a novel algorithm for discovering the nearest balanced states for any graph size in a fraction of the time. Our approach converges to the global optimum for the small graphs that the state-of-the-art binary linear programming (BLP) model computes. BLP does not work for graphs larger than 100,000 vertices while graphBpp seamlessly scales with the graph size to discover one or more nearest balanced states for the network. The state might not be optimal for a minimal number of edge switches, but it is close to optimal, and the algorithm produces a list of edges to switch to achieve the balanced state. We report the result on one computer for 1000, 2000, or 5000 iterations. In addition, we propose the use of gradient descent as a way to approximate the frustration index in linear time. We compared both graphBpp and graphL, and we deduce that the latter is efficient and generally yields more optimal balanced states as tested on Konect signed graphs.

### v3

@@ C.2 Implementation
The baseline implementation relies on the published Binary Linear Programming (BLP) code [Aref, 2021]. The binary linear model runs on a Jupyter Notebook in Python [Aref, 2021] and is based on a Gurobi mathematical solver and has several parameters like the termination parameter where we can set a time limit on how long the optimization process should last in Gurobi [Gurobi, 2008]. The binary terms in the objective function depend on the single AND constraints and two standard XOR constraints per edge, respectively. The graphBpp implementation extends the open-source implementation [Alabandi et al., 2021] to include and test proposed tree sampling strategies while keeping the original speedup optimization for finding fundamental cycles intact. The graphBpp algorithm works with different tree-sampling strategies. It is achieved using C++ and involves minimizing the number of loops used, incorporating OpenMP directives for parallel processing, and promptly freeing up memory resources when no longer needed. In the implementation of Algorithm 5, the code checks the total RAM size of the Linux system during runtime and the amount of memory currently used by the frustration cloud. These two values are compared in each iteration to decide how many balanced states can be stored. Two replacements in the ABS model’s objective function linearized two absolute value terms [Aref and Neal, 2021]. The code [Aref, 2021] was run with the following modifications: (1) -1 for the method parameter that indicates that the optimization method is automatic, and the setting will typically choose the non-deterministic concurrent method in the Gurobi’s documentation [Gurobi, 2008] for this linear programming problem; (2) the lazy parameter is set to 1 with enabled speedup; (3) thread parameter is set to multiprocessing.cpu_count(, and (4) the time limit for the model run is set to up to 30 hours. Note that the value of the lazy attribute influences how aggressively the model is constrained. A value of 1 allows the constraint to cut off a feasible solution. The code provided [Aref, 2021] generates random graphs based on the specified number of nodes, edges, and probability of negative edges. Our improvements to the code allow for the code to (1) accept the same input format as graphBpp and to (2) detect and eliminate duplicates, inconsistencies, self-loops, and invalid signs in the input graph. The code for graphBpp is available on GitHub, and the data it uses is publicly accessible. The references for these resources are [Anonimoys, 2023] for the code and [Leskovec and Krevl, 2014, Kunegis, 2013, He and McAuley, 2016] for the data. On the other hand, for the gradient-based heuristic, we set $\lambda$ to 1000 and the learning rate $\alpha$ to 0.001.
@@ Appendix D Tree-Sampling techniques for graphBpp
The Depth-First Search (DFS) algorithm [Cormen, 2009] is a graph traversal method characterized by its time complexity of O($|V|+|E|$). The traversal commences at a root vertex and continues to explore as deeply as possible along each pathway (branch) before backtracking. This process repeats until it reaches the vertex where all adjacent vertices have been visited. The Breadth first search (BFS) algorithm [Cormen, 2009] with time complexity O($|V|+|E|$) is a graph traversal approach in which the algorithm first passes through all vertices on the same level before moving on to the next level. A graph traversal technique is employed to visit all the vertices of a graph. A level is a group of equidistant vertices from the root vertex. We propose to use the randomized algorithms as follows: in each iteration, we shuffle and randomize a node’s neighborhood using a uniformly distributed random seed number before applying a static algorithm. The idea is that a vertex establishes a link to the first unvisited vertex based on the network’s randomized order of the adjacency list. Randomized Depth First Search (RDFS) algorithm transforms DFS into a non-deterministic algorithm by eliminating the static ordering of the adjacency lists. The time complexity of the DFS is known to be O($|V|+|E|$), where $|V|$ is the number of vertices and $|E|$ is the number of edges in the signed network. The algorithm also runs in linear time O($n$), where $n$ is the number of vertices adjacent to a specific vertex in the network, so the total time complexity is O($|V|+|E|$). Aldous-Broder algorithm with complexity O($|V|$) produces a random uniform spanning tree by performing a random walk on a finite graph with any initial vertex and stops after all vertices have been visited [Hu et al., 2021]. For the popular Kruskal’s algorithm [Wu et al., 2023] that has a time complexity O($|E|log|V|$) or O($|E|log|E|$), we intend to generate random spanning trees by assigning random weights to every edge in each iteration before running the algorithm. The method finds the minimum spanning tree of a connected and weighted graph. Randomizing the weights of Prim’s algorithm [Gass and Fu, 2013] [Wu et al., 2023] with complexity O($|V|^{2}$) can also generate random spanning trees.
@@ Appendix E Scaling graphBpp for Large Signed Graphs
The graphB+ algorithm [Alabandi et al., 2021] efficiently discovers fundamental cycles for the spanning tree $T$ and computes vertex and edge labels with linear time complexity. The algorithm requires a linear amount of memory, and the running time for balancing a cycle is linear in the length of the cycle times the vertex degrees but independent of the size of the graph. The labels here are specific values assigned to the vertices and edges after sampling a spanning tree as a preliminary step before finding and traversing the fundamental cycles and balancing them. These labels aid in finding these cycles and ensure efficient traversal. The algorithm speeds up the balancing to more than 14 million identified, traversed, and balanced fundamental cycles. Next, the algorithm traverses or visits each cycle in a specific order and notes which edge signs switched within each cycle to obtain an even number of negative edges along that cycle. The graphBpp, which builds on graphB+, scales the computation for large graphs as in Algorithm 5. First, we compute the number of balanced states that we can keep in the memory as $\mathcal{F}_{max}$ (Algorithm 5, line 3). Next, we keep only the $\mathcal{F}_{max}$ best-balanced states in $\mathcal{B}$, their count through all $k$ iterations in $\mathcal{C}$, and the number of switched edges for each of them in $\mathcal{S}(i)$. Note that there can be different balanced states of $\Sigma$ with the same number of switched edges. Finally, we compute the frustration index as a minimum of $\mathcal{S})$. The proposed approach scales well with the size of the signed graph, as we limit the number of the nearest balanced states that we keep in the memory based on their frustration index and capacity of the frustration cloud map in-memory storage, as illustrated in Algorithm 5. The comparison of balanced states now has up to $k$ iterations times $\mathcal{F}_{max}$ closest balanced states. We determine $\mathcal{F}_{max}$ so that the size of the frustration cloud in memory is smaller than $CAP$. In experiments, we define $CAP$ as 75% of the total RAM size, assuming that some vital system processes are running in the background. Note that $\mathcal{F}_{max}$ can be determined in the implementation by checking the current memory usage after every balanced state insertion and verifying that this insertion did not reach the limit. The benefit of saving balanced states in memory is that it allows us to record and study each state’s frequency and use them for other tasks, such as community detection and finding the largest balanced sub-graph.
@@
The size of the frustration cloud grows with the number of iterations as the probability of the previously unseen nearest balanced state grows. The size of the frustration cloud can become an issue for graphs with millions of vertices and vertices as the frustration cloud is too big for the main memory. The underlying data structure for the implementation in C++ is $std::map<std::string,std::pair<int,int>>$. The key, as discussed above, is of type string that represents a balanced state, and the pair stores two integers, one for the number of switches and the other for the frequency of the corresponding stable state. When comparing keys, the string keys are slower than integer keys in the map data structure. However, in our case, storing the balanced states using strings or integers should have a similar comparison performance, and the integer key approach would be much more complicated. First, if we use $std::map<int,std::pair<int,int>>$, then it would be difficult and more complex to represent the signed network using one integer key, whereas the string can intuitively concatenate all the edges with their signs to represent that graph. There is no direct way of representing this collection of edges forming the signed graphs using solely the integer key. If we were to use an integer as a key for the map, we would still have to loop through the edges in that key for the comparisons, and the cost would be $O(|E|)$ equivalent to using the string key.
@@
Our approach is different than the work by Tang et al. [Tang and Zhu, 2023] in several ways. First, we estimate the different types of latent variables, and we only use a random float vector $\Gamma$ of size equal to the number of vertices to determine the optimal membership of each vertex in the Harary subsets. Second, we are not modeling any signed networks. Our focus is to flip the sign of the edges after estimating the latent variables of each vertex and approximate the frustration index. Tang et al. did not explicitly and directly modify the edge signs of the graph. Third, we propose using a loss function to compute the frustration index and the gradients directly. Fourth, we use the vanilla gradient descent approach instead of the projected gradient descent to minimize our loss function. We simply threshold the latent variables after optimization (ex, if an element in the latent vector is 25 after the gradient-descent step, which is above 0, we assign that element to be 1). Fifth, our loss function is adopted from Du et al. [Du et al., 2016] in which they propose to measure imbalance for fully signed networks. The loss function used by Tang et al. is $L=\sum_{i<j}^{n}{|A_{ij}|}\frac{1+A_{ij}}{2}\eta_{ij}+|A_{ij}|\log(1-\Sigma(\eta_{ij})))$ where $\eta_{ij}=v_{i}v_{j}$, $A$ is the adjacency matrix, $\sigma$ is the sigmoid function, and $n$ is the number of nodes.
@@ G.2 The graphL Complexity Analysis
For every gradient update and computation, it is sufficient to loop over every edge in the signed graph to update the elements in the $\Gamma$ vector. The $\lambda$ is the number of gradient updates, and the total time complexity becomes O($\lambda*|E|$). We utilize the compressed sparse row (CSR) format to model the signed graph instead of the adjacency matrix because CSR scales better memory-wise. The construction of the adjacency matrix has a space complexity of O($V^{2}$), which isn’t computationally feasible when dealing with large signed graphs. Computing the imbTheccurs in constant time in each gradient update iteration does not affect the total time complexity.
@@ Appendix H More Results on the Comparison between BLP and graphBpp
The execution time of BLP and graphBpp with all seven tree-sampling techniques is reported on the log 10 scale in seconds in Figure 5 (bottom) as BLP takes 30 hours for larger datasets (far right navy bar in Fig. 5 (bottom)). RDFS-BFS is competitive with BFS in terms of frustration index as a percentage of the total number of edges in the graph (green and blue bars in Figure 5 (top)) with the small timing overhead for large graphs (Fig. 5 (bottom)): BFS produces 117,587 frustrations while BFS-RDFS produces 115,932 frustrations in 1000 iterations for the slashdot dataset. The Prim approach is too slow for large datasets, and the baseline BLP takes too long, or it does not complete. We tabulated the timing per iteration for each tree-sampling technique in Table 8. Since the time complexity of Prim is O($V^{2}$), the number of iterations is set to 1, and it was very inefficient and slow for large graphs such as WikiConflict. Moreover, Aldous-Broder (AB) did not terminate for DutchCollege because, in uncommon scenarios, AB would get stuck looping when performing a random walk after all the current vertex’s neighbors have been visited.
@@ Appendix I Concluding Remarks
There is more than one way to achieve balance in the network. The frustration index characterizes the optimal nearest balanced state where the minimum edge switches are required to achieve balance in the network. The tree-spanning approach to graph balancing produces the nearest balanced states, e.g., there can be no other balanced state nearest balanced state derived from [Rusnak and Tešić, 2021]. In this paper, we extend our findings and propose a novel algorithm for discovering the nearest balanced states for any graph size in a fraction of the time. Our approach converges to the global optimum for the small graphs that the state-of-the-art binary linear programming (BLP) model computes. BLP does not work for graphs larger than 100,000 vertices while graphBpp seamlessly scales with the graph size to discover one or more nearest balanced states for the network. The state might not be optimal for a minimal number of edge switches, but it is close to optimal, and the algorithm produces a list of edges to switch to achieve the balanced state. We report the result on one computer for 1000, 2000, or 5000 iterations. In addition, we propose the use of gradient descent as a way to approximate the frustration index in linear time. We compared both graphBpp and graphL, deducing that the latter is efficient and generally yields more optimal balanced states as tested on Konect signed graphs.